OpenStudy (jellybot23):

I'm so confused Transform this to rectangular form rcsc(θ)=5

satellite73 (satellite73):

rewrite as \[\frac{r}{\sin(\theta)}=5\] as a first step

satellite73 (satellite73):

then \[r=5\sin(\theta)\] now replace \(\sin(\theta)\) by \(\frac{y}{r}\)

satellite73 (satellite73):

if i lost you yet, let me know only two more steps

OpenStudy (jellybot23):

This is my first time trying these so they are quite confusing to me hehe. But I am pretty sure I am following. You are replacing sin(theta) with y/r because of the equation y=rsin(theta) right?

satellite73 (satellite73):

rigth

satellite73 (satellite73):

so now we are at \[r=\frac{y}{r}\]right ?

satellite73 (satellite73):

oh no wrong \[r=\frac{5y}{r}\] forgot the 5

OpenStudy (jellybot23):

Okay, yes, I see :)

jimthompson5910 (jim_thompson5910):

Here's how I did it. See the attached image. Either way, you'll get the same answer.

satellite73 (satellite73):

yeah same thing next would be \[r^2=5y\] and \[x^2+y^2=5y\]

OpenStudy (jellybot23):

Thank you so much both of you!! Can you help me with one more @satellite73

satellite73 (satellite73):

sure why not (if i can)

OpenStudy (jellybot23):

Haha! \[r=(4/2\cos \theta-3\sin \theta)\]

satellite73 (satellite73):

btw it might have been easier to replace \(\sin(\theta)\) byu \(\frac{y}{r}\)right away to get \[\frac{r^2}{y}=5\]in one step so you have lots of choices

satellite73 (satellite73):

\[r=\frac{4}{2\cos(\theta)-3\sin(\theta)}\]

OpenStudy (jellybot23):

Yes, that's it

satellite73 (satellite73):

i think it is a line in any case put \(\cos(\theta)=\frac{x}{r}\) and \(\sin(\theta)=\frac{y}{r}\) and do a bunch of agebra

satellite73 (satellite73):

\[r=\frac{4}{\frac{2x}{r}-\frac{3y}{r}}\]

OpenStudy (jellybot23):

That makes sense!

satellite73 (satellite73):

i would actually flip that sucker to make it less painful \[\frac{1}{r}=\frac{\frac{2x}{r}-\frac{3y}{r}}{4}\]

satellite73 (satellite73):

now multiply both sides by \(r\) you ca see they will go bye bye

OpenStudy (jellybot23):

Oh! Wow, that's amazing haha! And then just multiply by 4?

satellite73 (satellite73):

yeah i guess

satellite73 (satellite73):

you get something like \[4=2x-3y\] right?

OpenStudy (jellybot23):

That's what I got :)

OpenStudy (jellybot23):

Thank you so so much!

satellite73 (satellite73):

yw